*****PLEASE NUMBER AND BOLD ALL ANSWERS****

1. The customer service desk at Joe's Discount Electronics store receives 5 customers per hour on average. On average, each customer requires 10 minutes for service. The customer service desk is staffed by a single person. If modeled as an M/M/1 queuing system, what is the arrival rate λ and service rate  μ?

2. Continued to problem at Joe’s Discount Electronics store, what is the average waiting time in the entire system?

3.Continued to problem at Joe’s Discount Electronics store, What is the average number of customers in the entire system?
Answer #1

1.

Arrival rate (lambda) = 5 customers per hour

Service rate (mu) = 60/10 = 6 customers per hour

2.

Average waiting time in the system is given by the formula

W = 1/(mu-lambda) = 1/(6-5) = 1 hour

3.

Average number of customer in the system is given by the formula

L = lambda/(mu-lambda) = 5/(6-5) = 5 customers